MsBrains (ms-brains):

Can somebody please check my work? I need a second opinion to make sure it's correct. :) I will medal

OpenStudy (anonymous):

Sure, post it.

MsBrains (ms-brains):

I didn't show the work, but if that is a problem It will take me longer to copy it from my paper on here.

Atsie (atsie):

Its right :)

OpenStudy (anonymous):

Correct. \(5^2+4=29\) and \(6\times5-1=29\)

MsBrains (ms-brains):

I have 4 practice problems that need checking. :)

OpenStudy (anonymous):

Incorrect. \(5\times(-6)<6\times(-6)+6\) so \(30 < 30\) which isn't true.

Atsie (atsie):

From the way I see it, m > - 6. After rechecking it again

OpenStudy (anonymous):

To solve this, simplify the inequality. \(5m<6m+6\) so \(-m<6\) so \(m > -6\). THe only answer satisfying that is -5.

MsBrains (ms-brains):

Yeah, That is what I got too m > -6 How can it be -5?

OpenStudy (anonymous):

Because \(-5>-6\)

OpenStudy (mathmale):

Please note: You can check your own answers in questions of this type by substituting your "solution" into the original equation. if that equation come out "true," your "solution" is correct. Try to avoid dependence on other people to check your results.

MsBrains (ms-brains):

Alright. Thank you. Next one..

OpenStudy (anonymous):

Correct.

Atsie (atsie):

if it helps any, the alternate solutions for problem number two are m + 6 > 0 and 5m < 6 (m + 1) and I'm going to stay away from number three because i myself am unsure of that one. Sorry about that.

MsBrains (ms-brains):

Last one And I am not sure on this. I think it is B..

OpenStudy (anonymous):

\[5x-15y=20\]\[5x=20+15y\]\[5x-20=15y\]\[\frac{5}{15}-\frac{20}{15}=y\] Can you simplify this and see which answer it is?

OpenStudy (anonymous):

That last one should be 5x/15. Sorry.

MsBrains (ms-brains):

x-4/3

OpenStudy (mathmale):

If you know you can check your own work, then why are you still asking others to check your work for you? What will you do during a test, when you have no one to do that?

OpenStudy (anonymous):

Try again\[y=\frac{5x}{15}-\frac{20}{15}\]

MsBrains (ms-brains):

I have no idea. Everything I solve it to is incorrect. ;-;

OpenStudy (anonymous):

How can you simplify \(\frac{5}{15}\) and \(\frac{20}{15}\)?

MsBrains (ms-brains):

Simplify 5/15 to 1/3 and work from there..

MsBrains (ms-brains):

5/15 - 20/15 would = -1

MsBrains (ms-brains):

@tom982 You there?

MsBrains (ms-brains):

Is it A...?

OpenStudy (anonymous):

Close, but you can't just lose the x, \(y=\frac{5}{15}x-\frac{20}{15}=\frac{1}{3}x-\frac{4}{3}\)

MsBrains (ms-brains):

It's C!

MsBrains (ms-brains):

y=1/3 x - 4/3